# SE-VLN: A Self-Evolving Vision-Language Navigation Framework Based on Multimodal Large Language Models

### v2

@@ Introduction
Vision-language navigation (VLN), as a key technology connecting human natural language and robot visual navigation, aims to enable embodied agents to autonomously plan paths and complete visual navigation tasks in unseen environments based on human language instructions (Anderson et al. 2018). Despite its rapid development (Chen et al. 2024b; Schumann et al. 2024; Li et al. 2024; Lin et al. 2024), especially with the help of emerging LLMs (Zhou, Hong, and Wu 2024; Chen et al. 2024a; Li, Chen, and Lin 2024; Long et al. 2024), current VLN methods lack the advanced navigation capabilities of natural agents (e.g., bats, migratory birds, humpback whales) and shows critical shortcomings in autonomous evolution through experience. In fact, an experienced horse refines its neural circuitry through repeated journeys. Its navigation skill distills wisdom from experience and evolves autonomously through adaptation, providing a biological blueprint for existing VLN to overcome bottlenecks of ”data dependency” and ”scene generalization”.
Current VLN methods (Li, Chen, and Lin 2024; Zhou, Hong, and Wu 2024; Chen et al. 2024a) had significant limitations in terms of static experience utilization. They treated historical trajectory data as a fixed replay buffer, only using it to maintain decision consistency for the current task, and failed to extract reusable general knowledge from dynamic processes. Another drawback of existing VLN methods (Chen et al. 2024a; Zhou, Hong, and Wu 2024) lay in their limited reasoning capabilities. They relied on the static knowledge invocation of pre-trained models and failed to effectively integrate experience for multi-step decision-making. This stood in stark contrast to the biological intelligence of natural agents such as horses, which adjusted navigation paths by combining environmental memories with real-time perceptions, helping gain precise alignment between visual perceptions and language instructions when processing complex commands. Furthermore, the ”evolution” processes of existing VLN agents (Anderson et al. 2018; Lin et al. 2024, 2025) ware highly dependent on manual hyperparameter tuning and model iteration, which was completely different from the mechanism which biological agents used to achieve self-evolution through natural selection and neural plasticity.
@@
Researchers had extensively explored VLN methods that integrated memory and reasoning mechanisms to enhance the efficiency and robustness of agent navigation (Zhou, Hong, and Wu 2024; Chen et al. 2024b; Schumann et al. 2024; Li et al. 2024). The memory module played a crucial role in VLN agents, enabling them to store and review historical information such as explored paths (Chen et al. 2024a), identified landmarks (Zhan et al. 2024), or past observation data (Li, Chen, and Lin 2024). This capability was vital for handling long-range navigation tasks and making informed decisions in non-Markovian environments, as agents could utilize memory to optimize path selection and more effectively understand the connections between the current environment and historical states. Meanwhile, the reasoning module endowed agents with the ability to deeply understand language instructions and analyze environmental visual information. Agents could decompose abstract natural language instructions into a series of executable sub-goals and perform semantic analysis on visual inputs to identify relevant objects, comprehend spatial relationships, and infer feasible paths (Chen et al. 2024a; Zhan et al. 2024). This strong reasoning capability allowed agents to make rational decisions even when faced with ambiguous or incomplete instructions, and to extract key clues from visual scenes that guided navigation. However, these methods heavily relied on the inherent knowledge bases and reasoning capabilities of LLMs to improve navigation performance.
@@ A Retrieval-Augmented Thought-Based Reasoning Module
=$\mathcal{L}=\{l_{j}\}_{j=1}^{m}=\mathcal{F}_{\text{MLLM}}(\mathcal{I}),\quad\text{where\;}l_{j}\in\mathcal{V}_{\text{landmark}}$ (5)
@@
[…]
Where ||||| | denoted the sentence concatenation operation, and $\text{SBERT}(\cdot)$ represented the Sentence-BERT (Reimers 2019) encoder.

### v3

@@ Introduction
Vision-language navigation (VLN), as a key technology connecting human natural language and robot visual navigation, aims to enable embodied agents to autonomously plan paths and complete visual navigation tasks in unseen environments based on human language instructions (Anderson et al. 2018). Despite its rapid development (Chen et al. 2024b; Schumann et al. 2024; Li et al. 2024a; Lin et al. 2024), especially with the help of emerging LLMs (Zhou, Hong, and Wu 2024; Chen et al. 2024a; Li, Chen, and Lin 2024; Long et al. 2024), current VLN methods lack the advanced navigation capabilities of natural agents (e.g., bats, migratory birds, humpback whales) and shows critical shortcomings in autonomous evolution through experience. In fact, an experienced horse refines its neural circuitry through repeated journeys. Its navigation skill distills wisdom from experience and evolves autonomously through adaptation, providing a biological blueprint for existing VLN to overcome bottlenecks of ”data dependency” and ”scene generalization”.
Current VLN methods (Li, Chen, and Lin 2024; Zhou, Hong, and Wu 2024; Chen et al. 2024a) had significant limitations in terms of static experience utilization. They treated historical trajectory data as a fixed replay buffer, only using it to maintain decision consistency for the current task, and failed to extract reusable general knowledge from dynamic processes. Another drawback of existing VLN methods (Chen et al. 2024a; Zhou, Hong, and Wu 2024) lay in their limited reasoning capabilities. They relied on the static knowledge invocation of pre-trained models and failed to effectively integrate experience for multi-step decision-making. This stood in stark contrast to the biological intelligence of natural agents such as horses, which adjusted navigation paths by combining environmental memories with real-time perceptions, helping gain precise alignment between visual perceptions and language instructions when processing complex commands. Furthermore, the ”evolution” processes of existing VLN agents ware highly dependent on manual hyperparameter tuning (Anderson et al. 2018; Lin et al. 2024) and model iteration (Li et al. 2024b; Lin et al. 2025), which was completely different from the mechanism which biological agents used to achieve self-evolution through natural selection and neural plasticity.
@@
Researchers had extensively explored VLN methods that integrated memory and reasoning mechanisms to enhance the efficiency and robustness of agent navigation (Zhou, Hong, and Wu 2024; Chen et al. 2024b; Schumann et al. 2024; Li et al. 2024a). The memory module played a crucial role in VLN agents, enabling them to store and review historical information such as explored paths (Chen et al. 2024a), identified landmarks (Zhan et al. 2024), or past observation data (Li, Chen, and Lin 2024). This capability was vital for handling long-range navigation tasks and making informed decisions in non-Markovian environments, as agents could utilize memory to optimize path selection and more effectively understand the connections between the current environment and historical states (Chen et al. 2024a; Zhan et al. 2024). Meanwhile, the reasoning module endowed agents with the ability to deeply understand language instructions and analyze environmental visual information. Agents could decompose abstract natural language instructions into a series of executable sub-goals and perform semantic analysis on visual inputs to identify relevant objects, comprehend spatial relationships, and infer feasible paths (Chen et al. 2024a; Zhan et al. 2024). This strong reasoning capability allowed agents to make rational decisions even when faced with ambiguous or incomplete instructions, and to extract key clues from visual scenes that guided navigation. However, these methods mostly relied on the inherent knowledge bases and reasoning capabilities of LLMs to improve navigation performance.
@@ A Retrieval-Augmented Thought-Based Reasoning Module
$\mathcal{L}=\{l_{j}\}_{j=1}^{m}=\mathcal{F}_{\text{MLLM}}(\mathcal{I}),\quad\text{where\;}l_{j}\in\mathcal{V}_{\text{landmark}}$ (5)
@@
$\mathbf{q}=\text{SBERT}\left(\underset{j=1}{\overset{m}{\parallel}}l_{j}\right)\in\mathbb{R}^{768}$ (6)
Where $||$ denoted the sentence concatenation operation, and $\text{SBERT}(\cdot)$ represented the Sentence-BERT (Reimers 2019) encoder.